• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name dhamilton

    Log entry time 08:11:00 on April12,2008

    Entry number 231557

    keyword=Owl Shift Summary

    Owl Shift Summary
    D. Hamilton
    K. Adrashev

    pi0 production kinematics D

    00:00 - shift started with 3 microamps on target during run 2819.

    00:40 - rebooted ROC5.

    01:10 - Beam current increased to 3.5 microamps. Deadtime = 13%.

    01:35 - MCC called. They had some problems with the RF communication systems.

    02:30 - beam reatored to hall. Instability in beam position caused BB HV trip.

    05:00 - rebooted ROC5.

    08:00 - shift ended with 3.5 microamps on target during 2858.

    Runs taken during the shift:
    2819 - 2858

    Estimated total number of events:
    34 M (17 microamp hours)

    Bad runs:
    2821, 2823, 2825, 2828, 2842